# Pointwise convergence of ergodic averages with Möbius weight

### v1

@@ Abstract.
converge to 00 pointwise almost everywhere, provided that either: (i) $k\leq 3$ and $P_{i}$ are linear; (ii) $k$ is arbitrary and $P_{i}$ have distinct degrees; (iii) $k$ and $P_{i}$ are arbitrary, but $N$ is restricted to a certain explicit sequence of natural numbers. Specialising to $P_{1}(y)=y,P_{2}(y)=2y$, our results solve a problem of Frantzikinakis. For the proofs we establish some quantitative generalised von Neumann theorems for polynomial configurations that are of independent interest.
@@ 1. Introduction
The convergence properties of these averages as $N\to\infty$ have been studied intensively. The question of $L^{2}(X)$ norm convergence of these averages was settled, after a series of partial results, by Host–Kra [16] and Leibman [20]. The question of pointwise convergence, however, remains open and is the subject of the celebrated Furstenberg–Bergelson–Leibman conjecture [2, Section 5.5]. Pointwise convergence has so far been established in two notable cases, namely the case where $k=1$ and the case where $k=2$ and $P_{1}$ is linear. The former follows from celebrated work of Bourgain [3], and the latter follows from another well known work of Bourgain [4] if $P_{2}$ is linear and from a recent breakthrough of Krause, Mirek and Tao [18] if $P_{2}$ has degree at least $2$.
Let $\mu$ be the Möbius function, defined as $\mu(n)=(-1)^{k}$ if $n$ is the product of $k$ distinct primes, and $\mu(n)=0$ if $n$ is divisible by the square of a prime. In this paper, we shall consider the Möbius-weighted polynomial multiple ergodic averages
@@
Based on the Möbius randomness principle (see [17, Section 13]), it is natural to conjecture the following.
@@
(1.2) $\lim_{N\to\infty}\frac{1}{N}\sum_{n\leq N}\mu(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
@@
The corresponding $L^{2}(X)$ norm convergence result was proven by Frantzikinakis and Host [8]. The only case we are aware of where Conjecture 1.1 was previously known is the case $k=1$ and $P_{1}$ linear, proven in [6, Proposition 3.1] (see also [8, Theorem C] for a generalisation of that result to other multiplicative functions).
Our first main theorem settles Conjecture 1.1 for $k=2$ in full, for $k=3$ if $P_{1},P_{2},P_{3}$ are linear, and for $k\geq 3$ if $P_{1},\ldots,P_{k}$ have distinct degrees. In fact, we get a polylogarithmic rate of converge to 00. These results go beyond what is currently known about the unweighted averages (1.1).
@@
Let $k\in\mathbb{N}$, and let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients. Suppose that one of the following holds:
  1. (i)
    $k\leq 3$ and $P_{1},\ldots,P_{k}$ are linear polynomials;
  2. (ii)
    $P_{1},\ldots,P_{k}$ have pairwise distinct degrees.
Let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$. Let $(X,\nu,T)$ be a measure-preserving system. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$ and $A\geq 0$, we have
@@
    Specialising to $k=2$ and $P_{1}(y)=y,P_{2}(y)=2y$ (and $A=0$, $q_{1}=q_{2}=\infty$), Theorem 1.2 settles the Möbius case of Problem 12 of Frantzikinakis’ open problems survey [7] [^1] Frantzikinakis also asked about the convergence of the same multiple ergodic averages weighted by any multiplicative function $f\colon\mathbb{N}\to\mathbb{C}$, taking values in the unit disc, which has convergent means in the sense that $\lim_{N\to\infty}\mathbb{E}_{n\leq N}f(an+b)$ exists for all $a,b\in\mathbb{N}$. This general question remains open; see Remark 6.2.. This problem was also stated by Frantzikinakis and Host in [8].
@@
    In case (i), we can allow iterates of three different commuting transformations in the result; see Theorem 1.5 below.
@@
    It is likely that, at least for $k=2$, the region of $(q_{1},\ldots,q_{k})$ in the theorem could be improved to $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}\leq 1+\delta_{d}$ for some $\delta_{d}>0$, hence “breaking duality” in this problem. This is thanks to the $L^{p}$-improving estimates of Lacey [19] (see also [5]) and Han–Kovač–Lacey–Madrid–Yang [15]. See also [18, Section 11]. We leave the details to the interested reader.
@@
Our second main theorem shows that Conjecture 1.1 holds in general along a universal subsequence of $N$ (in fact, any sequence that grows fast enough).
Theorem 1.3.
Let $(N_{j})_{j\in\mathbb{N}}$ be any sequence satisfying $N_{j}\gg\exp(\exp(\exp((\log j)^{2})))$.
Let $k\in\mathbb{N}$, and let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients. Let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$. Let $(X,\nu,T)$ be a measure-preserving system. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$, we have
$\lim_{j\to\infty}\frac{\exp((\log j)^{3/2})}{N_{j}}\sum_{n\leq N_{j}}\mu(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
for almost all $x\in X$.
We also show that Conjecture 1.1 holds if we replace the Cesàro average in (1.2) by a weaker double-logarithmic average. In the rest of the paper, for a function $F\colon[N]\to\mathbb{C}$ we define its double-logarithmic average up to $N$ as
(1.3) $\mathbb{E}_{n\leq N}^{\log\log}F(n)\coloneqq\left(\sum_{2\leq n\leq N}\frac{F(n)}{n\log n}\right)/\left(\sum_{2\leq n\leq N}\frac{1}{n\log n}\right).$
From partial summation one sees that, for any bounded function $F$, the existence of the limit $\lim_{N\to\infty}\frac{1}{N}\sum_{n\leq N}F(n)$ implies the existence of the limit $\lim_{N\to\infty}\mathbb{E}^{\log\log}_{n\leq N}F(n)$, but the converse is not true in general.
Theorem 1.4.
Let $k\in\mathbb{N}$, and let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients. Let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$. Let $(X,\nu,T)$ be a measure-preserving system. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$, we have
$\lim_{N\to\infty}(\log\log N)^{c}\,\mathbb{E}_{n\leq N}^{\log\log}\mu(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
for almost all $x\in X$, for any $c>0$ small enough in terms of $P_{1},\ldots,P_{k},q_{1},\ldots,q_{k}$.
@@ 1.1. Ergodic averages with commuting transformations
Let a probability space $(X,\nu)$ and invertible, commuting, measure-preserving maps $T_{1},\ldots,T_{k}\colon X\to X$ be given. For any polynomials $P_{1},\ldots,P_{k}$ with integer coefficients and $f_{1},\ldots,f_{k}\in L^{\infty}(X)$, one can consider the more general polynomial ergodic averages with commuting transformations
@@
The commuting case seems to be much more difficult, since in the unweighted case (1.4) pointwise convergence is not currently known even for $k=2$ and $P_{1},P_{2}$ linear (see Problem 19 of [7]). However, we mention that Walsh [28] proved $L^{2}(X)$ norm convergence of the averages (1.4) in a groundbreaking work, and the Furstenberg–Bergelson–Leibman conjecture states that pointwise convergence should hold even with commuting transformations.
Our next theorem states that for the Möbius averages (1.5) we have pointwise convergence in the case $k=3$ and $P_{1},P_{2},P_{3}$ linear.
@@
Let $(X,\nu)$ be a probability space and let $T_{1},T_{2},T_{3}\colon X\to X$ be invertible, commuting, measure-preserving maps. Let $1<q_{1},q_{2},q_{3}\leq\infty$ satisfy $\frac{1}{q_{1}}+\frac{1}{q_{2}}+\frac{1}{q_{3}}<1$, and let $f_{1}\in L^{q_{1}}(X),f_{2}\in L^{q_{2}}(X),f_{3}\in L^{q_{3}}(X)$. Then, for any $A\geq 0$, we have
$\lim_{N\to\infty}\frac{(\log N)^{A}}{N}\sum_{n\leq N}\mu(n)f_{1}(T_{1}^{n}x)f_{2}(T_{2}^{n}x)f_{3}(T^{n}x)=0$
@@
Naturally, this theorem implies Theorem 1.2(i) by restricting to $T_{1},T_{2},T_{3}$ being powers of the same transformation.
It seems likely that also Theorems 1.3 and 1.4 could be extended to the case of commuting transformations by extending Theorem 4.2 that goes into their proof to the case of multivariate functions (which could likely be done with a more complicated PET induction scheme). We leave the details of this extension to the interested reader.
@@ 1.2. Prime ergodic averages
The arguments presented in this paper are not limited to polynomial ergodic averages weighted by Möbius function, but instead apply to similar ergodic averages with any weight that satisfies certain Gowers uniformity assumptions as well as some upper bound assumptions that are easy to verify; see Theorems 6.16.46.6. In particular, thanks to the quantitative Gowers uniformity estimates for the von Mangoldt function in [22][26], these general theorems can be applied to reduce the problem of convergence of polynomial ergodic averages weighted by the primes to the problem of convergence of the same averages weighted by integers with no small prime factors, which is an easier problem (though still highly nontrivial). Pointwise convergence of polynomial multiple ergodic averages weighted by the primes will be studied in a future joint work with Krause, Mousavi and Tao.
@@ 1.3. Further applications of the proof method
A key ingredient in the proofs of our main theorems are some new polynomial generalised von Neumann theorems with quantitative dependencies that we establish in Section 4. These results are likely to have applications also to other problems than ergodic averages, in particular to quantitative asymptotics for the Möbius and von Mangoldt functions along polynomial progressions and to bounds for sets of integers lacking progressions of the form $x,x+P_{1}(p-1),\ldots x+P_{k}(p-1)$ with $p$ prime and with $P_{1},\ldots,P_{k}$ polynomials of distinct degrees with $P_{i}(0)=0$. Such applications will be investigated in future works.
@@ 2. Proof ideas
We now give an overview of the arguments used to prove Theorems 1.21.31.4, presenting the steps of the proof in a somewhat different order than in the actual proof and focusing on the case of functions $f_{i}\in L^{\infty}(X)$ for simplicity.
@@
Let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients and with highest degree $d$. The starting point of the proofs of our pointwise convergence results is a lacunary subsequence trick combined with the Borel–Cantelli lemma and Markov’s inequality, which reduce matters to obtaining strong quantitative pointwise bounds for the averages
(2.1) $\left|\frac{1}{N^{d+1}}\sum_{m\leq N^{d}}\sum_{n\leq N}\mu(n)\phi(T^{m}x)f_{1}(T^{m+P_{1}(n)}x)\cdots f_{k}(T^{m+P_{k}(n)}x)\right|$
for $\nu$-almost all $x$ and all $\phi\in L^{\infty}(X)$. This reduction is presented in Section 6. For proving Theorem 1.2, we need bounds that save a large power of $(\log N)^{-1}$, whereas for proving Theorem 1.3 we need bounds that save a small power of $(\log\log N)^{-1}$ (and these suffice also for Theorem 1.4).
@@
We then establish (in Section 4) polynomial generalised von Neumann theorems for counting operators of the form (2.1), which allow bounding the averages (2.1) in terms of the $U^{s}[N]$ Gowers norm of $\mu$ for some $s\in\mathbb{N}$, or some weaker norm such as the $u^{s}[N]$ norm (defined in Section 3). See in particular Theorems 4.1 and 4.2. Crucially, the bounds we obtain for (2.1) in terms of these norms of $\mu$ are quantitative with polynomial dependencies. These polynomial generalised von Neumann theorems are proven, depending on the case, either using the Peluse–Prendiville degree-lowering theory [23][24] or repeated use of van der Corput’s inequality coupled with the PET induction scheme.
After establishing the desired generalised von Neumann theorems, we can conclude the argument by applying the strongest known quantitative bounds for the $U^{s}[N]$ and $u^{s}[N]$ norms of the Möbius function (presented in Lemma 5.1). In the case of Theorem 1.2(ii), we use the fact that for any $s\in\mathbb{N}$ the $u^{s}[N]$ norm of $\mu$ has an arbitrary power of logarithm of cancellation (which follows from the Green–Tao Möbius and nilsequences result [13] but could also be proven by more classical means). In the case of Theorem 1.2(i), the proof is finished by using a recent quantitative $U^{4}[N]$ Gowers norm estimate for $\mu$ due to Leng [22] (which saves an arbitrary power of logarithm). For Theorems 1.3 and 1.4, in turn, we use a quantitative $U^{s}[N]$ Gowers norm estimate for $\mu$, due to Tao and the author [26] (which saves a small power of double-logarithm).

### v2

@@ Abstract.
converge to 00 pointwise almost everywhere. Specialising to $P_{1}(y)=y,P_{2}(y)=2y$, this solves a problem of Frantzikinakis. We also prove pointwise convergence for a more general class of multiplicative weights for multiple ergodic averages involving distinct degree polynomials. For the proofs we establish some quantitative generalised von Neumann theorems for polynomial configurations that are of independent interest.
@@ 1. Introduction
The convergence properties of these averages as $N\to\infty$ have been studied intensively. The question of their $L^{2}(X)$ norm convergence was settled by Host–Kra [20] and Leibman [24] after a series of substantial progress by several authors. The question of pointwise convergence, however, remains open and is the subject of the celebrated Furstenberg–Bergelson–Leibman conjecture [2, Section 5.5]. Pointwise convergence has so far been established in two notable cases, namely the case where $k=1$ and the case where $k=2$ and $P_{1}$ is linear. The former follows from celebrated work of Bourgain [3], and the latter follows from another well known work of Bourgain [4] if $P_{2}$ is linear and from a recent breakthrough of Krause, Mirek and Tao [22] if $P_{2}$ has degree at least $2$.
Let $\mu$ be the Möbius function, defined by $\mu(n)=(-1)^{k}$ if $n$ is the product of $k$ distinct primes and by $\mu(n)=0$ if $n$ is divisible by the square of a prime. In this paper, we shall consider the Möbius-weighted polynomial multiple ergodic averages
@@
Based on the Möbius randomness principle (see [21, Section 13]), it is natural to conjecture the following.
@@
$\lim_{N\to\infty}\frac{1}{N}\sum_{n\leq N}\mu(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
@@
The corresponding $L^{2}(X)$ norm convergence result was proven by Frantzikinakis and Host [10] (this could also be deduced from the proof of [5, Theorem 1.3], combined with [17, Theorem 1.1]). The only case we are aware of where Conjecture 1.1 was previously known is the case $k=1$, proven in [7, Theorem 2.2] (see also [8, Proposition 3.1] for the case of a linear polynomial and [10, Theorem C] for a generalisation of that result to other multiplicative functions).
Our first main theorem settles Conjecture 1.1 in full. In fact, somewhat unusually for ergodic theorems, we get a quantitative (polylogarithmic) rate of converge. This result goes beyond what is currently known about the unweighted averages (1.1).
@@
Let $k\in\mathbb{N}$, and let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients. Let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$. Let $(X,\nu,T)$ be a measure-preserving system. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$ and $A\geq 0$, we have
@@
    Specialising to $k=2$ and $P_{1}(y)=y,P_{2}(y)=2y$ (and $A=0$, $q_{1}=q_{2}=\infty$), Theorem 1.2 settles the Möbius case of Problem 12 of Frantzikinakis’ open problems survey [9] [^1] Frantzikinakis also asked about the convergence of the same multiple ergodic averages weighted by any multiplicative function $g\colon\mathbb{N}\to\mathbb{C}$, taking values in the unit disc, which has a mean value in every arithmetic progression. Theorem 1.6 below makes progress on this more general question.. This problem was also stated by Frantzikinakis and Host in [10].
@@
    In the case where the $P_{i}$ are linear, we can allow iterates of different commuting transformations in the result; see Theorem 1.3 below.
@@
    It is likely that, at least for $k=2$ and $P_{1}$ linear, the region of $(q_{1},\ldots,q_{k})$ in the theorem could be improved to $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}\leq 1+\delta_{d}$ for some $\delta_{d}>0$, hence “breaking duality” in this problem. This is thanks to the $L^{p}$-improving estimates of Lacey [23] (see also [6]) and Han–Kovač–Lacey–Madrid–Yang [19]. See also [22, Section 11]. We leave the details to the interested reader.
@@ 1.1. Ergodic averages with commuting transformations
Let a probability space $(X,\nu)$ and invertible, commuting, measure-preserving maps $T_{1},\ldots,T_{k}\colon X\to X$ be given. For any polynomials $P_{1},\ldots,P_{k}$ with integer coefficients and $f_{1},\ldots,f_{k}\in L^{\infty}(X)$, one can consider polynomial ergodic averages with commuting transformations
@@
The commuting case seems to be rather more difficult, since in the unweighted case (1.2) pointwise convergence is not currently known even for $k=2$ and $P_{1},P_{2}$ linear (see Problem 19 of [9]). However, we mention that Walsh [35] proved $L^{2}(X)$ norm convergence of the averages (1.2) in a groundbreaking work, and the Furstenberg–Bergelson–Leibman conjecture asserts that pointwise convergence should hold also with commuting transformations.
Our next theorem states that for the Möbius averages (1.3) we have pointwise convergence in the case where the $P_{i}$ are linear.
@@
Let $k\in\mathbb{N}$. Let $(X,\nu)$ be a probability space and let $T_{1},\ldots,T_{k}\colon X\to X$ be invertible, commuting, measure-preserving maps. Let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$, and let $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$. Then, for any $A\geq 0$, we have
$\lim_{N\to\infty}\frac{(\log N)^{A}}{N}\sum_{n\leq N}\mu(n)f_{1}(T_{1}^{n}x)\cdots f_{k}(T_{k}^{n}x)=0$
@@
Naturally, this theorem implies Theorem 1.2 for $P_{i}$ being linear by taking to $T_{1},\ldots,T_{k}$ being powers of the same transformation.
It seems likely that also the general case of Theorem 1.2 could be obtained for commuting transformations by extending Theorem 4.2, which goes into its proof, to multivariate functions (which could likely be done with a more complicated PET induction scheme). We leave the details to the interested reader.
1.2. Multiplicative weights
We also consider more general weighted polynomial multiple ergodic averages
(1.4) $\frac{1}{N}\sum_{n\leq N}g(n)f_{1}(T^{P_{1}(n)})\cdots f_{k}(T^{P_{k}(n)}x),\quad x\in X,$
with $g\colon\mathbb{N}\to\mathbb{C}$ a function. Already for $k=1$ and $P_{1}$ linear, a necessary condition for the pointwise convergence of these averages is that $g$ has convergent means, meaning that $\lim_{N\to\infty}\frac{1}{N}\sum_{n\leq N}g(an+b)$ exists for all $a,b\in\mathbb{N}$ (this is seen by taking $X$ to be a finite set).
In Theorem 7.1 we will show, assuming a mild growth condition on $g$, that good decay bounds on the Gowers $U^{s}[N]$ norms of $g$ (or for certain weaker $u^{s}[N]$ norms that depend on the maximal correlation with polynomial phases) imply the convergence of the averages (1.4) to 00. Such results should have applications also in cases where $g$ has no arithmetic structure (for example, for random weights); however, here we focus on applications with $g$ being multiplicative [^2] We say that $g\colon\mathbb{N}\to\mathbb{C}$ is multiplicative if $g(mn)=g(m)g(n)$ whenever $m,n$ are coprime.
Frantzikinakis [9] conjectured that if $g\colon\mathbb{N}\to\mathbb{C}$ is $1$-bounded, multiplicative and hasconvergent means, then we have the pointwise almost everywhere convergence of
$\frac{1}{N}\sum_{n\leq N}g(n)f_{1}(T^{n}x)f_{2}(T^{2n}x).$
There is an obvious extension of this conjecture to polynomial multiple ergodic averages.
Conjecture 1.4.
Let $k\in\mathbb{N}$, and let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients. Let $g\colon\mathbb{N}\to\mathbb{C}$ be a multiplicative function taking values in the unit disc and having convergent means. Let $(X,\nu,T)$ be a measure-preserving system. Then, for any $f_{1},\ldots,f_{k}\in L^{\infty}(X)$, the limit
$\lim_{N\to\infty}\frac{1}{N}\sum_{n\leq N}g(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)$
exists for almost all $x\in X$.
While we are not able to prove this statement in full, we can prove it for a natural class of multiplicative functions, namely those satisfying a Siegel–Walfisz assumption (stated below). Most practically occurring multiplicative functions of mean 00 satisfy this property, and this class arises naturally in several problems in analytic number theory, in particular in connection with the Bombieri–Vinogradov theorem (see e.g. [15]).
Definition 1.5.
We say that a function $g\colon\mathbb{N}\to\mathbb{C}$ satisfies the Siegel–Walfisz assumption if the following hold:
  1. (1)
    $g$ is divisor-bounded: for some $C\geq 0$, we have $|g(n)|\leq d(n)^{C}$ for all $n\in\mathbb{N}$, with $d(n)$ denoting the number of positive divisors of $n$.
  2. (2)
    For all $A>0$ and $N\geq 3$ we have
    $\max_{1\leq a\leq q\leq(\log N)^{A}}\left|\sum_{\begin{subarray}{c}n\leq N\\ n\equiv a\ (\mathrm{mod}\ q)\end{subarray}}g(n)\right|\ll_{A}\frac{N}{(\log N)^{A}}.$
Examples of multiplicative functions satisfying the Siegel–Walfisz assumption include $h(n)d(n)^{j}\chi(n)n^{it}$ for any integer $j\geq 0$, real $t$ and Dirichlet character $\chi$, where $h\colon\mathbb{N}\to\mathbb{C}$ is any bounded multiplicative function that is “pretending” to be the Möbius function in the sense that $\sum_{p}\frac{1+\textnormal{Re}(h(p))}{p}<\infty$. [^3] That these functions are examples can be verified by using Perron’s formula and standard estimates for Dirichlet $L$-functions close to the $1$-line.
We are now ready to state a result on the pointwise convergence of multiple ergodic averages with a multiplicative weight satisfying the Siegel–Walfisz assumption.
Theorem 1.6.
Let $k\in\mathbb{N}$, and let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients and with distinct degrees. Let $1<q_{1},\ldots,q_{k}\leq\infty$ satisfy $\frac{1}{q_{1}}+\cdots+\frac{1}{q_{k}}<1$. Let $(X,\nu,T)$ be a measure-preserving system. Let $g\colon\mathbb{N}\to\mathbb{C}$ be a multiplicative function satisfying the Siegel–Walfisz assumption. Then, for any $f_{1}\in L^{q_{1}}(X),\ldots,f_{k}\in L^{q_{k}}(X)$, we have
$\lim_{N\to\infty}\frac{1}{N}\sum_{n\leq N}g(n)f_{1}(T^{P_{1}(n)}x)\cdots f_{k}(T^{P_{k}(n)}x)=0$
for almost all $x\in X$.
Note that we allow the function $g$ to be unbounded, hence proving pointwise convergence in some cases not covered by Conjecture 1.4.
Somewhat curiously, Theorem 1.6 requires an argument that is rather different from our proof of Theorem 1.2; both proofs will be discussed in Section 2.
@@ 1.3. Prime ergodic averages
The arguments presented in this paper are not limited to polynomial ergodic averages weighted by the Möbius function, and indeed apply to similar ergodic averages with any weight that satisfies certain Gowers uniformity assumptions as well as some weak upper bound assumptions that are easy to verify; see Theorem 7.1. In particular, thanks to the quantitative Gowers uniformity estimates for the von Mangoldt function in [25], these general theorems can be applied to reduce the problem of convergence of polynomial ergodic averages weighted by the primes to the problem of convergence of the same averages weighted by integers with no small prime factors, which is an easier problem (though still highly nontrivial). Pointwise convergence of polynomial multiple ergodic averages weighted by the primes will be studied in a future joint work with Krause, Mousavi and Tao.
@@ 1.4. Further applications of the proof method
Key ingredients in the proofs of our main theorems are some new polynomial generalised von Neumann theorems with quantitative dependencies that we establish in Section 4. These results are likely to have applications also to other problems, such as to bounds for sets of integers lacking progressions of the form $x,x+P_{1}(p-1),\ldots x+P_{k}(p-1)$ with $p$ prime and with $P_{1},\ldots,P_{k}$ polynomials of distinct degrees with $P_{i}(0)=0$. Such applications will be investigated in future works.
@@ 2. Proof ideas
We now give an overview of the arguments used to prove Theorems 1.2 and 1.6, presenting the steps of the proof in a somewhat different order than in the actual proof and focusing on the case of functions $f_{i}\in L^{\infty}(X)$ for simplicity.
@@
We begin with Theorem 1.2. Let $P_{1},\ldots,P_{k}$ be polynomials with integer coefficients and with highest degree $d$. Thefirst step for the proofs of our pointwise convergence results is a lacunary subsequence trick, which combined with the Borel–Cantelli lemma and Markov’s inequality reduces matters to obtaining strong quantitative pointwise bounds of the form
[…]
for $\nu$-almost all $x$ and all $\phi\in L^{\infty}(X)$, with $A$ large enough. This reduction is presented in Section 7.
@@
We then establish a polynomial generalised von Neumann theorem for counting operators on the left-hand side of (2.1), which bounds the averages in (2.1) in terms of the $U^{s}[N]$ Gowers norm of $\mu$ for some $s\in\mathbb{N}$; see Theorem 4.2. This result is proven by repeated applications of van der Corput’s inequality coupled with the PET induction scheme. Crucially, the bounds we obtain for (2.1) in terms of the Gowers norm $\|\mu\|_{U^{s}[N]}$ are quantitative with polynomial (in fact, linear) dependencies.
After establishing this generalised von Neumann theorem, we can conclude the proof by applying the strongest known quantitative bounds for the $U^{s}[N]$ norms of the Möbius function (see Lemma 5.1), which save an arbitrary power of logarithm, thanks to recent work of Leng [25] that builds on the work of Leng–Sah–Sawhney [26].
In the case of Theorem 1.6, we repeat the lacunary subsequence argument to reduce to (2.1), with $\mu$ replaced by a multiplicative function $g$ satisfying the Siegel–Walfisz assumption. If one were now to apply Theorem 4.2 again, one would not be able to obtain a sufficiently strong bound on the $U^{s}[N]$ norm of $g$, since the Leng–Sah–Sawhney inverse theorem [26] is quasipolynomial rather than polynomial. We overcome this by establishing a different polynomial generalised von Neumann theorem, Theorem 4.1, that (perhaps unexpectedly at first) allows bounding (2.1) in terms of a weaker norm than the $U^{s}[N]$ norm of the weight function. This weaker norm, called the $u^{s}[N]$ norm and defined in (3.3), expresses the maximal correlation of the weight $g$ with a polynomial phase of degree at most $s-1$. The proof of Theorem 4.1 draws motivation from the Peluse–Prendiville degree lowering theory [28][29]. The proof proceeds by induction on the length of the progression and involves showing that the first two functions in a weighted progressions can be assumed to be “locally linear” phase functions in a suitable sense. This conclusion is then boosted to global linearity with some extra work, which allows reducing the length of the progression, hence completing the induction.
Since the $u^{s}[N]$ norm already involves correlations with polynomial phases, we are able to bypass the need for the inverse theorem for the $U^{s}[N]$ norm when working with distict degree polynomials. In Subsection 5.2, we show (using in particular a restriction to typical factorisations and bilinear estimates for polynomial exponential sums) that if $g$ is multiplicative and satisfies the Siegel–Walfisz assumption, then $g$ is close in $L^{1}[N]$ norm to a function $\widetilde{g}$ whose $u^{s}[N]$ norms decay faster than any power of logarithm. This together with Theorem 4.1 mentioned above suffices for concluding the proof of Theorem 1.6. The Siegel–Walfisz assumption gives just the right decay for (2.1): with any weaker assumption we would not be able to prove this (although the averages (1.4) should still converge).
We lastly remark that the approach based on Theorem 4.1 also gives a different and arguably simpler proof of Theorem 1.2 for distinct degree polynomials that is independent of any inverse theorems for the Gowers norms and only uses classical analytic number theory input (the only property needed of the Möbius function is an exponential sum estimate that essentially goes back to the work of Vinogradov [34] from 1939; see Remark 5.2).